# Conditional Dependence via U-Statistics Pruning

### v1

@@ VI Conclusions
The classical HSIC measure for marginal statistical dependence has been reinterpreted as a correlation measure on a finite but high-dimensional space based on windowed steering vectors. From the original mapping, we can establish an insightful connection with kernels by letting the dimension grow to infinity. Instead of being merely a reinterpretation of the HSIC, this vision demonstrates its main power under the more involved conditional dependence framework. In particular, the illustrated identification of the HSIC as a sample covariance estimator opens the possibility of leveraging U-Statistics for this task. Thanks to this formulation, we can provide a novel route for performing statistical conditioning by pruning data pairs based on the pairwise differences of the confounder under suspicion. Furthermore, the proposed measure of conditional dependence does not require matrix inversions, which has the advantage of reduced computational complexity and the avoidance of addressing ill-conditioned matrices. Further work should study the potential of the proposed method with richer data sets.

### v2

@@ VI Conclusions
This work has proposed a proof of concept for a modification on the classical HSIC under the conditional dependence framework, named C-HSIC. Its interpretation as a correlation measure on a finite but high-dimensional space allows an insightful connection with kernels by letting the dimension grow without bound. Moreover, it opens the possibility of leveraging U-Statistics for this task. Thanks to this formulation, we can provide a novel approach for performing statistical conditioning by pruning data pairs based on the pairwise differences of a potential confounder. Furthermore, the proposed measure of conditional dependence does not require matrix inversions, which has the advantage of reduced computational complexity and the avoidance of addressing ill-conditioned matrices. Numerical illustrations have shown that C-HSIC is capable of measuring both conditional dependence and independence in two different scenarios. Further research should study the potential of the proposed method with richer data sets, as well as provide a thorough comparison with other methods for measuring conditional dependence.